A) The half-life for the decay of fluorine-18 is 109.8 min. What percent of the initial amount of fluorine-18 will remain after 439.2 min? The decay of fluorine-18 is a first order reaction.

B) The decay of carbon-14 is used to date objects related to human civilization. The decay has a half-life of approximately 5730 years. The book that "nobody can read" or the Voyich manuscript was dated by faculty at the University of Arizona. It was found that 93.0% of the initial amount of carbon-14 reamined in 2011. Inwhat year was the Voyich manuscript produced?

ANSWER BOTH PARTS A AND B
Answer #1

A) The decay constant is calculated:

k = 0.693 / t1 / 2 = 0.693 / 109.8 min = 6.3x10 ^ -3 min-1

The percentage of decayed fluoride is calculated:

% F = (1 - exp-k * t) * 100 = (1 - exp -6.3x10 ^ -3 * 439.2) * 100 = 93.71%

B) The decay constant is calculated:

k = 0.693 / 5730 years = 1.2x10 ^ -4 years-1

You have the equation of previous decay:

7% = (1 - exp 1.2x10 ^ -4 * t) * 100

T = 605 years is cleared

It occurred in the year 2011 - 605 = 1406.

If you liked the answer, please rate it in a positive way, you would help me a lot, thank you.